OpenStudy (anonymous):

A local florist charges $8 for a red rose and $10 for a white rose. He sold twice as many red roses as white roses for a total income of $8,554 for the day. How many roses did he sell all together?

OpenStudy (anonymous):

Lets say he sold x number of white roses Then according to the question red roses are twice the number of the white roses sold which means 2x Total income earned will now be x + 2x = $8554 3x = $8554 => x = $2851 Which means he earned $2851 by selling white roses. But question is we need to find the number of white roses sold In the question they mentioned cost of each white rose is 10$ so number of white roses sold = 2851/10 = 285 :) And red roses are obviously twice the white roses which is equal to = 570 So total roses sold will be 570 + 285 = 855 :)